I have been trying these questions over 10 times i have failed this test 10 times i really need help to try and understand this please show me how to do this. The ordered pair (4,12) is a solution for which equation
a.)y=1/6x
b.)y=3x
c.)y=1/3x
d.)y-6x

Answer by richwmiller(17219)

The ordered pair mean (x, y) so x=4 and y=12 (4,12)
a)12=1/6*4
b)12=3*4
c)12=(1/3)*4
d)12=6*4
Which one comes out 12=12?